OpenStudy (anonymous):

Don't understand how to work out this problem, any help is greatly appreciated

OpenStudy (anonymous):

Here's the problem attached

OpenStudy (anonymous):

Ik it's small but just zoom in to see it

OpenStudy (anonymous):

do you know that the radius \(\perp\) the cord at its middle point?

OpenStudy (anonymous):

the radius is perpendicular to the cord at it's middle point? Ok so what exactly is the cord again?

OpenStudy (anonymous):



OpenStudy (anonymous):

hence, if AB =19.2 , then you can apply pythagorean to get x, right?

OpenStudy (anonymous):

would be 9.6^2+x^2=?

OpenStudy (anonymous):

=12^2

OpenStudy (anonymous):

how did you get 12^2?

OpenStudy (anonymous):

diameter =24 --> radius =12

OpenStudy (anonymous):



OpenStudy (anonymous):

Okay so it would be 9.6^2+x^2=12^2?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Cool Thanks mind helping with a few more?

OpenStudy (anonymous):

not sure but if i can, i will but Mersi here, she is the best in geometry, not worry :)

OpenStudy (anonymous):

thanks appreciate it

OpenStudy (anonymous):

so how did you get 19.2 when it only said 16 on that last problem?

OpenStudy (anonymous):

don't get what you mean

OpenStudy (anonymous):

19.2 is given, I don't get it from anywhere

OpenStudy (anonymous):

oh yeah my bad XD

OpenStudy (anonymous):

And what is the "last problem"? I saw just only 1 problem

OpenStudy (anonymous):

Here it is

OpenStudy (anonymous):

hey, it is exactly the same, just diameter =16 and the chord = 25.6

OpenStudy (anonymous):

so the diameter isn't longer than 16?

OpenStudy (anonymous):

nope, exactly 16 --> radius =8

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

awesome

OpenStudy (anonymous):

it can be radius =16, let me check

OpenStudy (anonymous):

Yes, in this case, it is radius =16

OpenStudy (anonymous):

So the diameter is 32 right

OpenStudy (anonymous):

how can we know? we know it because the radius is the hypotenuse in the right triangle, and it MUST be longer than the leg which is \(\dfrac{25.6}{2}=12.8\)

OpenStudy (anonymous):

okay I get it

OpenStudy (anonymous):

if it is 8, we don't have hypotenuse > the leg That is the way to check in the case they give us an ambiguous information

OpenStudy (anonymous):

Cool I don't wanna bug you so can you help with this 1 last problem?

OpenStudy (anonymous):

we have formula to find it out, the ratio : \(\dfrac{AB}{AD}=\dfrac{AC}{AE}\) only one term you don't know is AD, quite easy, right?